A man in a rowboat that is a = 4 miles from the nearest point A on a straight shoreline wishes to reach a house located at a point B that is b = 10 miles farther down the shoreline. He plans to row to a point P that is between A and B and is x miles from the house, and then he will walk the remainder of the distance.
Suppose he can row at a rate of 3 mi/hr and can walk at a rate of 5 mi/hr. If T is the total time required to reach the house, express T as a function of x.

Answer by Theo(13342)

my understanding of this problem is shown in the following diagram.

$$$

the rowboat is at point R.
the shorted distance point on the shore if point A.
the hours is at point B.

the distance from the rowboat to point A is 4 miles which is the length of line RA.

the distance from point A to the house at point B is 10 miles which is the length of line AB

the distance from point P to the house at point B is x miles which is the length of line PB.

this makes the distance from point A to point P equal to (10 - x).

the distance from the rowboat to point P is equal to the hypotenuse of the right triangle RAP which is the line RP.

that makes the distance equal to sqrt(4^2 + (10-x)^2).

simplify that to get distance from point A to P equal to sqrt(x^2-20x+116).

the basic formula to use is r * t = d
r is the rate.
t is the time.
d is the distance.

the man rows the boat at 3 miles per hour.
the man walks at 5 miles per hour.

let T1 equal the time it takes to get from point R to point P rowing the boat.
let D1 equal the distance from point R to point P.
you get r * t = d becomes 3 * T1 = D1
since D1 = sqrt(x^2-20x+116), you get 3 * T1 = sqrt(x^2-20x+116)
solve for T1 to get T1 = sqrt(x^2-20x+116)/3

let T2 equal the time it takes to get from poinnt P to point B walking.
let D2 equal the distance from point P to point B.
you get r * t = d becomes 5 * T2 = D2
since D2 = x, you get 5 * T2 = x
solve for T2 to get T2 = x/5

let the total time it takes to get from point R to point B through point P equal to T.
you get T = T1 + T2 = sqrt(x^2-20x+116)/3 + x/5

that's your solution.

you can combine that equation into one common denominator to get:

T = (5*sqrt(x^2-20x+116) + 3x)/15

the two equations of T, shown above, are equivalent.
they produce the same value for T.

the domain of the T function is 0 < x < 10, where x represents the distance from point P to point B.

x cannot be 0 because then point P would be the same point as point B.
x cannot be 10 because then point P would be the same point as point A.

that's because the problem statement says that point P is between point A and point B and can therefore not be at either of them.

either of the equations involving T will be your solution, depending on how you want to show them.